# Compactifications of phylogenetic systems and electrical networks

### v2

@@ Abstract.
We demonstrate a graphical map, a new correspondence between circular electrical networks and circular planar split systems. When restricted to the planar circular electrical case, this graphical map finds the split system associated uniquely to the Kalmanson resistance distance of the dual network, matching the induced split system familiar from phylogenetics. This is extended to compactifications of the respective spaces, taking cactus networks to the newly defined compactified split systems. The graphical map preserves both components and cactus structure, allowing an elegant enumeration of induced phylogenetic split systems.
@@ 1. Introduction
At first glance the scientific pursuits of phylogenetics and circuit design seem to have little in common. In one we seek to uncover ancestral connections in a mysterious past, in the other to build electrical connections that achieve a desired effect. However, in both cases we are trying to find a structure—a mathematical network—which fits a known list of parameters. The resistance matrix for a circular planar electrical network turns out to be identical to a specialized phylogenetic distance matrix. Thus methods from phylogenetics are immediately useful for circular planar electrical networks.
We leverage several tools for the study of networks that are usually seen in differing contexts: Kron reduction and circular minors on the one hand, split networks on the other. The former is named for Gabriel Kron, who is responsible for much of the early work connecting electrical circuits, linear algebra, and graph theory. The Kron reduction [12] refers to a replacement of a circular electrical network with a less interconnected version which is indistinguishable from the original, from the point of view of its terminal, or boundary, nodes. The response matrix of the original is indeed the Laplacian of the Kron reduced network.
@@
Spaces of circular electrical networks have been studied thoroughly over the last couple of decades. Collected early results are in the book by Curtis and Morrow [7]. More recently, the state of the art for circular planar electrical networks is represented by Kenyon and Wilson [28, 29, 30], Kenyon and Hersh [25], and Gao, Lam, and Xu [24]. The space of circular planar electrical networks is an embedded slice of the totally nonnegative Grassmannian, and thus projected into the amplituhedron, as shown in several papers by Galashin, Karp, and Lam [17, 18, 19, 20, 33]. Connections to plabic graphs (which represent classes of directed networks) have been made by Galashin, Postnikov, and Williams [21]. Recognizing the planarity of a network and recovering its conductances both saw improvements in efficiency [1, 30].
@@
Spaces of circular split systems are better known via phylogenetics, as in the book by Steel [36], and in recent papers by Catanzaro, Gambette, Balvočiūtė, Bryant, and Spillner [6, 22, 2]. Devadoss and Petti [11] showed that the space of circular split systems is an intuitive extension of the Billera-Holmes-Vogtmann space of phylogenetic trees [3], exhibiting a natural projection into the compactification of the real moduli space of curves [8, 9]. It is studied as a simplicial complex by Terhorst [37], and is related to the BME polytope and STSP polytope by Devadoss, Forcey, Scalzi, and Durell [10, 13, 16].
@@
In our previous works [11, 15, 14], we study the space of circular split systems and a unique correspondence between these and circular electrical networks. We review these results in Section 2. We describe convenient maps between circular electrical networks and circular split systems in Section 3. The latter are simpler combinatorial structures well-studied in the field of phylogenetics, so our maps can be seen as invariants of the electrical networks. Section 4 introduces compactified split systems and extends the correspondence to electrical cactus networks. In particular, our graphical map $\mathbf{\xi}$ takes any compactified circular network (cactus network, planar or not) to a compactified planar split system. Our main result follows, a summary of Theorems 8 and 13 in Section 4:
@@
For a planar cactus network $N$, the map $\mathbf{\xi}$ coincides with both the Kalmanson $\mathbf{\rho}$ and the induced $\mathbf{\sigma}$ split systems of the planar dual $N^{*}$. That is, $\mathbf{\xi}(N)=\mathbf{\rho}(N^{*})=\mathbf{\sigma}(N^{*})$ and $\mathbf{\xi}^{\prime}(N^{*})=\mathbf{\rho}(N)=\mathbf{\sigma}(N)$.
Moreover, if the network $N$ itself is planar, then $\mathbf{\xi}(N)$ gives a shortcut to the alternative, well-known ways to find the associated split system via the Kalmanson map. Indeed, we show that the split systems in the image of this multi-named map have a particular form, known as faithful [15] — those systems which are made up of splits displayed by some network. These faithful systems are an important subcomplex of the space of circular systems.
Finally, Section 5 proves enumeration results (summarized by Table 1, page 1) across the spectrum of spaces and maps between them (showcased by Figure 6, page 6). Theorem 1 has an immediate application for counting the faithful split systems, resulting in the following (expanding Corollary 22) :
@@
The number of cells in the complex of faithful split systems with $n$ boundary nodes is
@@
where $j_{0},\dots,j_{n}$ are ordered non-negative integers, $\mathcal{P}_{j_{i}}$ is the set of Ptolemy diagrams (from [26]) with $j_{i}$ vertices, and $t(s)$ is the number of optional edges of $s$.
We obtain this by observing that including the induced system of a network with multiple connected components requires extension to the compactification of the space of systems, and vice versa: the induced systems with multiple components arise from cactus networks in the compactification of the space of electrical networks. This presents a challenge to our counting methods, but our new graphical maps solve the problem by undoing that switch: they preserve both the connected components and the cactus structure.
@@ 2.1. Circular Electrical Networks
Physically, a network $N$ is made of conducting Ohmic wires with $n$ exposed terminals that can be tested in order: We apply unit voltage to each of the terminals in turn while grounding all the remaining $n-1$ terminals. This defines the response matrix $M(N)$, recording the current at terminal $j$ of $N$ when the voltage is applied to terminal $i$ as entry $M_{ij}$. Indeed, two circular electrical networks are electrically equivalent if they have the same response matrix. An equivalence class is planar if any representative is planar. The Kron reduction $K(N)$ is an invariant of the equivalence class of $N$, and several other invariants are discussed in [7].
A circular electrical network is a graph with a selected set of $n$ terminals (or boundary nodes) labeled by $[n]$ and arranged on a circle. The rest of the graph lies inside the circle, and the interior nodes are unlabeled. The edges are weighted with positive real numbers which typically stand for the conductance of each connection. A circular planar electrical network $N$ is one that can be represented with its terminals on a bounding circle and with no crossed edges in the disk. Usually the terminals are assumed to be numbered in clockwise order, and throughout this paper planarity is with respect to that order. [^1] We will consider planarity with respect to any circular order in a future work.
Figure 1(a) shows an example of a circular planar network $N$, where the non-labeled edges have conductance 1. This is a running example, and we show how to calculate the rest of Figure 1 in the following sections. The dual $N^{*}$ shown in part (d) is calculated in Figure 5. The conductances of the Kron reduction of the dual are given by the off-diagonal entries of the response matrix $M(N^{*})$, where
$M(N^{*})\ =\ \begin{bmatrix}-23/14&11/28&0&0&1/28&3/14&0&1\\ 11/28&-183/56&1&1&19/56&15/28&0&0\\ 0&1&-1&0&0&0&0&0\\ 0&1&0&-1&0&0&0&0\\ 1/28&19/56&0&0&-39/56&9/28&0&0\\ 3/14&15/28&0&0&9/28&-29/14&1&0\\ 0&0&0&0&0&1&-2&1\\ 1&0&0&0&0&0&1&-2\\ \end{bmatrix}$
@@
More practically, we could use an ohmmeter to test the resistance (impedance) between pairs of our terminals. We record these results as the resistance matrix $W(N)$, where $W_{ij}$ is the effective resistance between $i$ and $j.$ The entries $W_{ij}$ are a metric on the terminal nodes, as shown in [32]. Figure 2(a) shows a nonplanar case. Here are $M(N)$ and $W(N)$ for Figure 2(a):
@@
Figure 2(b) displays the Kron reduction for the network in part (a). Note that $N$ and $K(N)$ have the same response matrix, and are therefore equivalent as circular electrical networks. In fact, the response matrix $M(N)$ is the Laplacian of the Kron reduction $K(N)$, and thus $K(N)$ can be viewed as a visualization associated with $M(N)$. Curtis and Morrow [7] defined a space of response matrices $\mathbf{\Omega}_{n}$ as follows:
@@
The space $\mathbf{\Omega}_{n}$ is stratified into cells as a CW-complex. We use the same notation for the space and the face poset of cells. [^2] The poset of cells is denoted $EP_{n}$ in [1] For instance the enumeration of these cells $a_{n}=|\mathbf{\Omega}_{n}|$ is given recursively by
@@
This formula is shown in [1] and gives the OEIS sequence [A111088]. Each cell of dimension $k$ in $\mathbf{\Omega}_{n}$ corresponds to an (unweighted) electrical equivalence class of networks, and can be minimally represented by choosing conductance weights for a planar non-unique critical (or reduced) network with $k$ edges. A cell $x$ is contained in another cell $y$ when the network corresponding to $x$ can be found by either deleting or contracting edges of $y$. The left side of Figure 3 shows critical networks for each of the 52 cells of $\mathbf{\Omega}_{4}$. Each column corresponds to cells of $\mathbf{\Omega}_{4}$ of a fixed dimension with $f$-vector $\langle 1,6,14,16,10,4,1\rangle$.
Figure 3. Combinatorial action of $\mathbf{\xi}$ on the cells of the space $\mathbf{\Omega}_{4}$ to $\mathbf{\Psi}_{4}$.

### v3

@@ Abstract.
We describe a graphical map: a visual and numerical correspondence between circular electrical networks and circular planar split systems. When restricted to the planar circular electrical case, this graphical map finds the split system uniquely associated with the Kalmanson resistance distance of the dual network, matching the induced split system familiar from phylogenetics. This correspondence is extended to compactifications of the respective spaces, taking cactus networks to the cactus split systems defined herein. The graphical map preserves both network components and cactus structure, allowing an elegant enumeration of induced phylogenetic split systems. We introduce the global spaces of circular planar electrical networks and circular split systems. These are CW complexes, and the 0-cells of each are counted by the Bell numbers as opposed to the Catalan numbers.
@@ 1. Introduction
1.1. Motivation
Recently it was demonstrated that a favorite tool of phylogenetics researchers, the split system, is guaranteed to exhibit the large scale structure of a simple electrical resistor network [13]. This highlights a mathematical analogy between two problems: reconstructing a circuit in a ‘black box’ from surface measurements, and rebuilding a genetic history from existing population data.
Figure 1. Example labels for the cells in the spaces we consider, for $n=4$: (a) circular planar electrical networks, (b) cactus networks (c) circular split systems and (d) cactus circular split systems . The total numbers of these cells are listed in Table 1, row 4, and the entire sets for (a) and (c) are shown in Figure 4.
In this paper we carry the analogy further. A printed electrical circuit is made of conductors in a network, with terminal nodes on the exterior available for connections. Some of those exterior nodes might be duplicates, short-circuited together so that they are interchangeable in terms of their use. Similarly, monozygotic twins can be seen as genetically the same for the purpose of reconstructing a family tree from the DNA samples of living individuals. On a larger scale, basic DNA sequences of extant individuals from a biologically diverse ecosystem can be equated when the species match. Allowing nodes, either terminals or taxa, to be identical in the limit of being infinitely close to each other corresponds to taking an unbounded space and compactifying it. Here we define and relate the compactification of phylogenetic split systems (from the DNA examples) to the well-studied compactification of circular planar electrical network space.
$n$ $|EP_{n}|$ $|{\mathbf{\Omega}}_{n}|=|P_{n}|$ $|\mathbf{\Psi}_{n}|$ $|\mathbf{\Psi}_{n}|$ $|\mathbf{\xi}(EP_{n})|$ $|\mathbf{\sigma}(\mathbf{\Omega}_{n})|=|\mathbf{\xi}({\mathbf{\Omega}}_{n})|$
nodes cells of $\mathbf{\Omega}_{n}$ cells of $E_{n}$ cells of $\mathbf{\xi}({\mathbf{\Omega}}_{n})$ cells of $\mathbf{\xi}({E}_{n})$
1 1 1 1 1 1 1
2 2 3 2 3 2 3
3 8 15 8 15 8 15
4 52 105 64 117 49 102
5 464 945 1024 1565 373 839
OEIS [A111088] [A001147] [A006125] [A136654] [?] [?]
Table 1. Spaces: numbers of cells in each CW complex. Left to right these are circular planar electrical networks, cactus planar electrical networks, circular split systems, and cactus circular split systems. The last two columns count the cells in the range of our maps from electrical networks to split systems. Examples are in Figure 1.
From a geometric combinatorial perspective, our interest lies in comparing the spaces of electrical networks and split systems as CW-complexes with analogous cell-decompositions. We will review the definitions in the following sections, but for now we show some examples of each, to introduce our notation. $\mathbf{\Omega}_{n}$ is the space of circular planar electrical networks with $n$ exterior nodes, in notation from [7]. $\mathbf{\Omega}_{n}$ has its poset of cells denoted $EP_{n}$ in [1]. The cells correspond to equivalence classes of critical networks such as (a) in Figure 1. Subfigure 1 (c) shows an example of a split system in $\mathbf{\Psi}_{4}.$ The compactification of $\mathbf{\Omega}_{n}$ is the space of cactus networks, denoted $E_{n}$ in [33]. The cells of $E_{n}$ we denote as $\mathbf{\Omega}_{n},$ with an example in (b) of Figure 1. The compactification of $\mathbf{\Psi}_{n}$ is defined in this paper and is denoted $\mathbf{\Psi}_{n}$, with an example in (d). In the interest of readability, we often use one name for both the CW-complex and the poset of cells: we will refer to $|\mathbf{\Omega}_{n}|$ as the number of cells in $\mathbf{\Omega}_{n}$ and to $|\mathbf{\Psi}_{n}|$ as the number of cells in $\mathbf{\Psi}_{n}.$ Similarly, we will use $\mathbf{\Omega}_{n}$ interchangeably with $E_{n}.$
Further, we want to consider the sub-complexes of split systems that are the range of the graphical map $\mathbf{\xi}$ which we describe and relate to the Kalmanson and induced phylogenetic mappings. In Figure 1 (c) is the image of (a) under $\mathbf{\xi}$, and (d) is the image of (b). There are fewer cells in that range than split systems in general: the implied mapping on cells is not onto. Neither is it an injection of cells. However the graphical map does preserve the connected components and cactus structures, and this provides us a way to count cells in the range either in either case, whether we want just the ordinary split systems or their compactifications as well. Table 1 shows the results of counting the cells of all dimension in the various complexes.
@@ 1.1. Motivation
Spaces of circular electrical networks, both planar and not planar, have been studied thoroughly over the last couple of decades. Collected early results are in the book by Curtis and Morrow [7]. More recently, the state of the art for circular planar electrical networks is represented by Kenyon and Wilson [28, 29, 30], Kenyon and Hersh [25], and Gao, Lam and Xu [23]. The compactified space of circular planar electrical networks is an embedded slice of the totally nonnegative Grassmannian, and thus projected into the amplituhedron, as shown in several papers by Galashin, Karp, and Lam [16, 34, 17, 18, 19]. Connections to plabic graphs (which represent classes of directed networks) have been made by Galashin, Postnikov, and Williams [20]. Recognizing the planarity of a network and recovering its conductances both saw recent improvements in efficiency [1, 30].
@@
Spaces of circular split systems are better known via phylogenetics, as in the book by Steele [38], and in recent papers by Catanzaro, Gambette, Balvočiūtė, Bryant, and Spillner [6, 21, 2]. Devadoss and Petti [11] showed that the space of circular split systems is an intuitive extension of the Billera-Holmes-Vogtmann space of phylogenetic trees [3], exhibiting a natural projection into the compactification of the real moduli space of curves [8, 9]. It is studied as a simplicial complex by Terhorst [39], and is related to the BME polytope and STSP polytope by Devadoss, Forcey, Scalzi, and Durell [10, 12, 15].
In our recent papers [13, 14], we demonstrated an injective mapping from electrical equivalence classes of circular planar electrical networks to circular split systems. Every planar response matrix is a Kalmanson metric, and thus a unique weighted circular split system. To see if a given circular split system is the image of an electrical network is more complicated. For a given Kalmanson metric one can first reorder the matrix to match any cyclic order respected by its circular split system. Then there is a unique associated response matrix, which can be checked for nonnegative circular minors to decide if it is a planar electrical network. If so, then we show in [13] that the circular planar electrical network has the same overall tree structure as the split system network of the resistance matrix. Then the algorithm from Chapter 9 of [7] can reconstruct a critical subnetwork for any part not made of bridges, using the strand diagram. In [24] the authors show how to check for planarity directly using the (reordered) resistance metric. This is done via the Lam embedding into the nonnegative Grassmannian, but with a variation that takes the resistance metric as input. Then in [24] the authors show an alternative reconstruction of the electrical network again using the resistance metric directly to find the strand matching.
Our map $\mathbf{\xi}$ takes circular planar electrical cactus networks to cactus circular split systems.The map is quite simple in its operation on matrices: it is the identity map. This is true for the cactus networks as well, although in that case the map takes a set of matrices to the same set of matrices. The identity map is clearly a homeomorphism. Thus the facts about the space $E_{n}$ of cactus networks as shown in [33] and [17] are also true of the image of $\mathbf{\xi}$ in the compactified space of circular split systems. It is topologically a ball, and isomorphic to a certain subspace of the nonnegative Grassmannian. The map $\mathbf{\xi}$ often takes several cells of $E_{n}$ and embeds those cells inside of a single cell of the space of cactus circular split systems. However, the map does respect the cell structure in that, for cells, $x\leq y$ in the domain iff $\mathbf{\xi}(x)\leq\mathbf{\xi}(y)$ in the range. Thus it induces a new regular CW-complex structure on the circular electrical networks. Those cells in $E_{n}$ are labeled by their Kron reductions. Careful study of these cells in $E_{n}$ is important future work.
In this paper we first introduce cactus split systems and extend the maps to the domain of electrical cactus networks. Then we expand to global spaces: while traditionally the objects have a given cyclic order of $[n]$ we instead allow any cyclic order, and identify appropriate equivalent networks. That is how classical phylogentic spaces are studied, both the space of trees $BHV_{n}$ and the space of split networks (with all trivial splits) $CSN_{n}.$ The combinatorics become much more complex in the global case. With the equivalence in place even the basic enumeration of the objects becomes difficult. That is not surprising when we see how they relate to famous hard problems, specifically facets of the Symmetric Travelling Salesman polytope. Table 2 lists the total numbers of cells in our new global spaces, both compact and non-compact. The last column there lists the number of cells in the image of the global cactus networks, under our map $\mathbf{\sigma}$ that takes a cactus network to displayed splits on the parts of the partition into its connected components.
$n$ $|\mathbf{\Omega}^{global}_{n}|$ $|\mathbf{\Omega}^{global}_{n}|$ $|\mathbf{\Psi}^{global}_{n}|$ $|\mathbf{\Psi}^{global}_{n}|$ $|\mathbf{\sigma}(\mathbf{\Omega}^{global}_{n})|$
1 1 1 1 1 1
2 2 3 2 3 3
3 8 15 8 15 15
4 70 133 112 169 124
5 1466 2397 6976 ? ?
OEIS [?] [?] [?] [?] [?]
Table 2. Global spaces: numbers of cells.
We have the following applications in mind:
  1. (1)
    How can a network be constructed (more quickly) to meet specified measurements of electrical properties on its boundary?
  2. (2)
    How can we determine (more quickly) from those measurements whether a network is planar?
  3. (3)
    How many ways can a given circular planar electrical network be reordered at its boundary while retaining its characteristic measurements and planarity?
  4. (4)
    Can those potential reorganizations be obtained from the network’s electrical properties, measured at the boundary?
We insert the parenthetical “more quickly” since well-known methods based on positivity answer the first two questions, as in [7]. The application of such methods and our improvements is clear: a circular planar electrical network with desired responses at its terminal nodes is a circuit that can be printed on a board with terminal pins. The ability to reorder those terminal nodes while retaining the character and planarity is important to integrating the circuit into a larger design. Placing component circuits into a large array, with terminals available to connect to each other, requires first choosing the convenient ordering of the terminals on each circuit. Combining these queries, we can also make specific requests: if a circuit is planar, but we wish to add or cross wires, can the resulting circuit be made planar again?
1.2. Results
@@
We describe convenient maps between circular electrical networks and circular split systems. The latter are simpler combinatorial structures well-studied in the field of phylogenetics, so our maps can be seen as invariants of the electrical networks. Our graphical map $\mathbf{\xi}$ (defined in Section 3) takes any compactified circular network (cactus network, planar or not) to a cactus planar split system. The latter are defined in Section 4. We show how to count them, with Corollary 18: A formula for the the number $|\mathbf{\Psi}_{n}|$ of unweighted cactus split systems for a given cyclic order on $[n]$ is as follows:
$|\mathbf{\Psi}_{n}|=\frac{1}{n+1}\left(\sum_{j_{0}+\dots+j_{n}=n}\left(\prod_{i=0}^{n}|\mathbf{\Psi}_{j_{i}}|\right)\right),$
where the sum is over ordered lists of non-negative integers summing to $n$, and $|\mathbf{\Psi}_{n}|$ is the number of unweighted (non-compact) circular split systems: $|\mathbf{\Psi}_{n}|=2^{n\choose 2}.$
Some resulting numbers are shown in Table 1. The corollary follows via Lagrange inversion from the recursive functional equation of species composition, $A=B\circ(X\cdot A)$, as shown in Section 5. The general principle here is that the compactification $A$ of a space of networks is related to the original space of ordinary networks $B$ by that functional equation. Thus this corollary also applies to the counting of equivalence classes of electrical cactus networks: the number $|\mathbf{\Omega}_{n}|$ of cells for cactus networks can be found by:
where $|\mathbf{\Omega}_{n}|$ is the number of distinct classes of non-compact electrical networks. For instance, (5(52)+10(8)2+10(2)3+10(4)+5(1))/5=(2(4)1)!!.55210821023104515double-factorial241(5(52)+10(8)2+10(2)3+10(4)+5(1))/5=(2(4)-1)!!.( 5 ( 52 ) + 10 ( 8 ) 2 + 10 ( 2 ) 3 + 10 ( 4 ) + 5 ( 1 ) ) / 5 = ( 2 ( 4 ) - 1 ) !! . The double factorial formula was shown in [33], but the larger formula is new. [^1] Our formula is implied by an inverse formula given by Paul D. Hanna in entry [A111088] of [37].
A third instance of this principle is forthcoming. First though, if the network $N$ itself is planar, then $\mathbf{\xi}(N)$ is a shortcut to the alternative, well-known ways to find the associated split system:
@@
For a planar cactus network $N$, the map $\mathbf{\xi}$ coincides with both the Kalmanson and the induced split systems of the planar dual. That is, $\mathbf{\xi}(N)=\mathbf{\rho}(N^{*})=\mathbf{\sigma}(N^{*}).$ Respectively, we have $\mathbf{\rho}(N)=\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*}).$
Theorem 1 is a summary of Theorems 8 and 13 in Section 4, as extended to the compactified case via Theorem 16. Indeed, we show that the split systems in the image of this multi-named map have a particular form, known as faithful [14] — those systems which are made up of splits displayed by some network. These faithful systems are an important subcomplex of the space of circular systems. Theorem 1 has an immediate application for counting these faithful split systems. If the original network has multiple components, the faithful (induced) split system will be a cactus network, and vice versa. Thus counting the induced systems is more difficult. On the other hand, the graphical system has the same number of components and cactus form as the original system. Thus we can use our recursive counting principle to relate numbers of split systems for these faithful forms in the compact and non-compact cases, in Corollary 23 by the now recognizable formula:
Section 5 proves enumeration results (summarized by Table 1, page 1) for our new bands in the spectrum of spaces (showcased by Figure 7, page 7). We define a class of plabic tilings for sets of polygons and then we finish with the following (expanding Corollary 23) :
@@
The number of cells in the compactification of the complex of faithful split systems with $n$ boundary nodes is
@@
where $j_{0},\dots,j_{n}$ are ordered non-negative integers, ${\mathbf{\mathcal{T}}}_{j_{i}}$ is the set of plabic tilings of a polygon with $j_{i}$ vertices, and $t(s)$ is the number of boundary edges of the shaded regions of $s$.
Finally, in Section 6 we introduce the globalization of the classical spaces and begin their study. We point out some initial results, such as the fact that the global spaces allow all cyclic orders, so their compactifications allow all partitions rather than just non-crossing partitions of $[n].$ Specifically, this means that the 0-cells of the CW-complex in each global case will be counted by the Bell numbers; the 0-cells in the non-global case are enumerated by the Catalan numbers. Section 7 collects some corollaries that relate our theorems to the global cases. We count the number of consistent cyclic orders for a network and describe obstructions to a given order being circular.
@@ 2.1. Circular Electrical Networks
Physically, a general electrical network $N$ is made of conducting Ohmic wires with $n$ exposed terminals that can be tested in order: We apply unit voltage to each of the terminals in turn while grounding all the remaining $n-1$ terminals. This defines the response matrix $M(N)$, recording the current at terminal $j$ of $N$ when the voltage is applied to terminal $i$ as entry $M_{ij}$. Indeed, two circular electrical networks are electrically equivalent if they have the same response matrix. An equivalence class is planar if any representative is planar.
A circular electrical network is a graph with its selected set of $n$ terminals (or boundary nodes) labeled by $[n]$ and arranged on a circle. The rest of the graph lies inside the circle, and the interior nodes are unlabeled. The edges are weighted with positive real numbers which typically stand for the conductance of each connection. A circular planar electrical network $N$ is one that can be represented with its terminals on a bounding circle and with no crossed edges in the disk. Usually the terminals are assumed to be numbered in clockwise order. We will consider planarity with respect to any cyclic order in Section 6.
@@
The Kron reduction $K(N)$ is an invariant of the equivalence class of $N$. Recall that equivalent networks are defined by having the same response matrix. For planar networks, planar representations of two in the same class are related by a sequence of moves including replacing a series of edges, replacing parallel edges, deleting superfluous edges, and the $Y-\Delta$ move. Those moves and several other invariants are discussed in [7]. The latter include the set of connections across the network, and the associated set of positive circular minors calculated from the response matrix—all of which are nonnegative for planar networks. Another invariant is the sub-collection of networks of a given equivalence class with the minimal number of edges, called the critical networks. We won’t need most of these invariants for this paper, but they are implicitly counted when we count the cells of the CW-complexes. Each cell is made of response matrices that all share the values of these (unweighted) electrical invariants. The weighted split system associated to a network is an electrical invariant as well, as shown in [13]. The unweighted set of splits is also an invariant of the electrical equivalence class, but like the unweighted Kron reduction it is strictly weaker than the set of connections or the set of unweighted critical networks.
Figure 2(a) shows an example of a circular planar network $N$, where the non-labeled edges have conductance 1. This is a running example, and we show how to calculate the rest of Figure 2 in the following sections. The dual $N^{*}$ shown in part (d) is calculated in Figure 6. The conductances of the Kron reduction of the dual are given by the off-diagonal entries of the response matrix $M(N^{*})$, where this matrix is computed as the Schur complement of the graph Laplacian with respect to the interior nodes (we show examples in [13]). Here we only show the weight of one edge $M(N^{*})_{2,6}=15/28.$
More practically, we could use an ohmmeter to test the resistance (impedance) between pairs of our terminals. We record these results as the resistance matrix $W(N)$, where $W_{ij}$ is the effective resistance between $i$ and $j.$ The entries $W_{ij}$ are a metric on the terminal nodes, as shown in [32].
Figure 3(a) shows a nonplanar case. Here are $M(N)$ and $W(N)$ for Figure 3(a):
Figure 3(b) displays the Kron reduction for the network in part (a). Note that $N$ and $K(N)$ have the same response matrix, and are therefore equivalent as circular electrical networks. In fact, the response matrix $M(N)$ is the Laplacian of the Kron reduction $K(N)$, and thus $K(N)$ can be viewed as a visualization associated with $M(N)$. The rest of Figure 3 will be explained in the section on maps.
Curtis and Morrow [7] defined a space of response matrices $\mathbf{\Omega}_{n}$ as follows:
@@
The space $\mathbf{\Omega}_{n}$ is of is stratified into cells as a CW-complex. The poset of cells is denoted $EP_{n}$ in [1]. We often use the same notation for the space and the face poset of cells, determined by context. For instance the enumeration of these cells $a_{n}=|\mathbf{\Omega}_{n}|=|EP_{n}|$ is given recursively in [1] (and gives the OEIS sequence [A111088]):
@@
The space $\mathbf{\Omega}_{n}$ has dimension (${n\choose 2}.$ Each element in a cell of dimension $k$ in $\mathbf{\Omega}_{n}$ can be minimally represented by choosing positive conductance weights for a (non-unique) planar critical (or reduced) network with $k$ edges. A cell $x$ is contained in another cell $y$ when the network corresponding to $x$ can be found by either deleting or contracting edges of $y$. The left side of Figure 4 shows critical networks for each of the 52 cells of $\mathbf{\Omega}_{4}$. Each column corresponds to cells of $\mathbf{\Omega}_{4}$ of a fixed dimension with $f$-vector $(1,6,14,16,10,4,1)$.
Figure 4. Cells of the space $\mathbf{\Omega}_{4}$ on the left, $\mathbf{\Psi}_{4}$ on right. The image $\mathbf{\xi}(\mathbf{\Omega}_{4})\subset\mathbf{\Psi}_{4}$ is inside the dashed line. The highlighted system on lower right is the image of the 4 networks highlighted at the top on the left.